# Exploring Influence Factors on LLM Suitability for No-Code Development of End User IoT Applications

### v1

@@ 5.3. Implications for Other Research Fields
This work focuses on understanding the performance of LLMs in generating IoT automation software for non-technical users. Beyond the question of how well LLMs can perform these tasks and what causes their behavior, introducing LLM-based automation for the IoT raises interesting research questions outside the domain of computer systems research.
For example, our proposal will require additional research in human-computer interaction. The usability and accessibility of LLMs in the domain of IoT automation should be investigated further, focusing not just on the performance of LLMs but also investigating the perception that users have of the system. It is an open question whether an unsuccessful request leads to frustration with the system or more interest and interaction as users are given an opportunity to refine their prompts. Similarly, the cognitive load on users as they formulate their intentions as well as their trust in the system could be quantified.
@@ 6. Conclusion
LLM-supported no-code platforms leverage the natural language understanding capabilities of LLMs to generate functional code based on user inputs, enabling seamless customization without requiring any technical expertise. Understanding the factors influencing the performance of LLM-powered no-code platforms can give insights on how to optimize the real-world development process, ensuring reliability, optimizing efficiency and enhancing the user experience.
In this paper, we aim to find the influencing factors that affect the performance of LLM-powered no-code IoT platforms, considering of LLM choice, prompt language variance, and LLM linguistic background. We provide valuable insights into the design and optimization of future platforms. Specifically, model selection has the most significant impact on performance. A general-purpose LLM with advanced natural language understanding, which prioritizes providing concrete results over comprehensive ones, is generally preferred. A LLM which the linguistic background aligns with input language also showcase an outperforming performance. Additionally, the influence of prompt language varies across different LLMs and task complexity levels, which should be carefully considered in the design of no-code platforms. Specifically, for LLM with advanced natural language understanding and multilingual capabilities, e.g., GPT-4o, the original user prompts should be kept to ensure the best performance. While for LLMs with either in lightweight or with limit in natural language understanding capability, i.e., GPT-4o-mini and Copilot, adding a translation step can improve the performance.
@@ Acknowledgements.
Partially funded by the Sponsor Bundesministerium für Bildung und Forschung (BMBF, German Federal Ministry of Education and Research) https://www.bmbf.de/bmbf/en in the scope of the Software Campus 3.0 (Technische Universität Berlin) program – Grant #01IS23068.

### v2

@@ 5.5. Implications for Other Research Fields
This work focuses on understanding the performance of LLMs in generating software for non-technical users. Beyond the question of how well LLMs can perform these tasks and what causes their behavior, introducing LLM-based automation for user-oriented platforms raises interesting research questions outside the domain of computer systems research.
For example, our proposal will require additional research in human-computer interaction. The usability and accessibility of LLMs in the domain of end-user oriented automation should be investigated further, focusing not just on the performance of LLMs but also investigating the perception that users have of the system. It is an open question whether an unsuccessful request leads to frustration with the system or more interest and interaction as users are given an opportunity to refine their prompts. Similarly, the cognitive load on users as they formulate their intentions as well as their trust in the system could be quantified.
@@ 6. Conclusion
LLM-supported NCDPs leverage the natural language understanding capabilities of LLMs to generate functional code based on user inputs, enabling seamless customization without requiring any technical expertise. Understanding the factors influencing the performance of LLM-powered NCDPs can give insights on how to optimize the real-world development process, ensuring reliability, optimizing efficiency and enhancing the user experience.
In this paper, we aim to find the influencing factors that affect the performance of LLM-powered NCDPs, considering of LLM choice, prompt language variance, LLM linguistic training background, and an error-informed few-shot setting. We provide valuable insights into the design and optimization of future platforms. Specifically, model selection has the most significant impact on performance. A general-purpose LLM with advanced natural language understanding, which prioritizes providing concrete results over comprehensive ones, is generally preferred. A LLM which the linguistic background aligns with input language also showcase an outperforming performance. Additionally, the influence of prompt language varies across different LLMs and task complexity levels, which should be carefully considered in the design of NCDPs. Specifically, for LLM with advanced natural language understanding and multilingual capabilities, e.g., GPT-4o, the original user prompts should be kept to ensure the best performance. While for LLMs with either in lightweight or with limit in natural language understanding capability, i.e., GPT-4o-mini and Copilot, adding a translation step can improve the performance. Furthermore, incorporating an error-informed few-shot approach can improve LLM performance in NCDPs by providing task-specific feedback, particularly for coding-oriented and medium-performing models. However, its effect remains secondary to model choice, and practical use requires weighing the additional engineering effort and resource costs against expected benefits.
@@ Acknowledgements.
Partially funded by the Sponsor Bundesministerium für Forschung, Technologie und Raumfahrt (BMFTR, German Federal Ministry of Research, Technology and Space) https://www.bmftr.bund.de/EN/Home/home_node.html in the scope of the Software Campus 3.0 (Technische Universität Berlin) program – Grant #01IS23068.